# Training Machine Learning models at the Edge: A Survey

### 05 Mar 2024 v1

@@ VI-B Reinforcement learning
Reinforcement learning (RL) has been successfully applied in the past on different problems in areas such as robotics, recommendation systems, video games and automatic vehicles [226, 227], making RL a promising and interesting direction for edge learning. However, The training of RL models in resources constrained environments is often limited by high compute and memory requirements from gradient computations [228], making the application of RL in the context of edge learning challenging. Despite these challenges, multiple RL approaches have been proposed for training ML models in the edge. Among them, Federated Reinforcement Learning (FRL) [229, 230] is a promising approach that allow multiple RL agents to learn optimal control policies for a series of devices with slightly different dynamics [231]. FRL is employed to achieve diverse objectives including personalization [232, 233], IoT traffic management [234, 235], Autonomous Systems [231] and resource allocation for unmanned aerial vehicle (UAV) [236]. FedMC [95] integrates RL models trained by multiple edge devices into a general model based on a meta-learning approach. FedGPO is an RL-based aggregation technique for FL introduced in [54], and aims to optimizes the energy-efficiency of FL while guaranteeing model convergence. On the other hand, [237] introduces DDQN-Trust, a trust-based double deep Q-learning-based selection algorithm for FL that takes into account the trust scores and energy levels of the IoT devices to make appropriate scheduling decisions and integrate it with the main FL aggregation techniques (FedAvg, FedProx, FedShare and FedSGD). Other Federated Reinforcement Learning methods include [75, 238, 239, 240, 241, 242]
@@ VI-C Semi-supervised learning
As mentioned in the section VI-A using unlabeled data to train on the edge is very promising considering the vast amount of unlabeled data in edge and end devices [32]. Semi-supervised learning, constituting a machine learning paradigm, seeks to leverage both labeled and unlabeled data for specific learning tasks [249], presenting a means to harness the vast amount of unlabeled data at the edge, circumventing the challenges associated with supervised and unsupervised learning. Semi-supervised learning also provides other advantage on the edge, as training datasets are often not completely defined before training and might need to be completed using real-time data [250]. Naturally, several methodologies have surface to adapt semi-supervised learning on the edge, using federated learning based approaches [251, 252], or other learning techniques [253, 254, 255, 211, 256, 250, 250, 257].
@@ VII-B Smart Technologies
Smart cities are urban ecosystems designed using IoT technologies to solve urban life problems and improve residents quality of life [279]. In this context, edge learning has been proposed to solve different challenges, mainly for its ability to leverage ML capabilities while preserving network bandwidth and reducing the charge on cloud servers. In [280], a cloud-aided edge learning based on knowledge fusion for smart lighting system has been proposed. Another application of ML in the edge is in smart grid systems where ML is needed to improve demand forecasting and automated demand response, as well as to analyse data related to energy use and obtain energy consumption patterns [281], detect anomalies [282], improve communications [283] and security [284] in the system. Other applications in smart cities include the detection of abnormal and dangerous activities [285, 286], pedestrian detection [287], water consumption forcasting [42, 288] and reducing congestions in intelligent traffic systems [289, 290]. Other more general EL approaches for mart cities include [291, 292, 293, 276, 252].
@@ VII-C Autonomous vehicles
Autonomous vehicles are vehicles that can operate without human intervention, they utilize sensor technologies, AI, and networking to navigate and make decisions [295]. Autonomous vehicles include self-driving cars trucks, buses, drones, unmanned aerial vehicles (UAVs), and even small robots. As demonstrated in [296] offloading deep learning tasks to edge devices or servers can improve the inference accuracy while meeting the latency constraint, which makes edge learning prefectly suitable for this use-cases, and as expected there has been extensive research done in this area.
Unmanned Aerial Vehicles (UAVs) are by far the most prominent use of EL for autonomous vehicles. In [297], a Synchronous Federated Learning (SFL) structure for multi-UAVs was proposed. That aims to resolve device privacy concerns that comes from sending raw data to UAV servers, as well as UAVs’ limited processing or communication resources. On the other hand, [298] propose a model-aided federated MARL algorithm to coordinate multiple UAVs on data harvesting missions with limited knowledge about the environment, significantly reducing the real-world training data demand. As mentioned previously in section VII-B, [277] aims to assist farming opperations using UAVs that measure soil moisture and ambient environmental conditions and [299] proposes a model to derive computation specifications for learning-based visual odometry (VO) from physical characteristics of UAVs. Other EL applications for UAVs include [236, 300, 301, 302, 303, 299, 304, 305, 306, 307]
@@ VII-D Recommendation systems and personalization
Personalized crowdsourced livecast are another part of personalization methods that might benefit from being offloded to the edge. In [327] the rapid development of crowdsourced livecast and the challenges in providing personalized Quality of Experience (QoE) to viewers is discussed, and it introduces an intelligent edge-learning-based framework called ELCast, which integrates Convolutional Neural Networks (CNN) and deep reinforcement learning (DRL) in edge computing architectures for personalized crowdcast recommendation. In the area related to video game, personalisation involves constructing a system capable of adapting video game rules and content to better suit some aspect of the player preferences, personality, experience and performances [328]. Although not yet explored in edge and on-device learning, [329] propose a Deep Q network model to personalize games based on user-interaction on the edge.
@@ VII-E Human Activity Recognition
Human activity recognition (HAR) refers to the automation of the identification and categorization of the various activities performed by humans and their interactions with the environment [330]. Since personalization for HAR has been shown to improved the results and performances of these systems [331], training the model on the edge using Incremental learning or meta learning approaches can help achieve that while increasing the privacy and reduce the bandwidth consumption. PILOTE [127] proposes an incremental learning based approach for HAR that is intended to be trained on edge devices with extremely limited resources and provides reliable performance in handling catastrophic forgetting. While [332] proposes an approach that proposes a personalizable lightweight CNN model for HAR. As well as a training algorithm to find personalization-friendly parameters, for boosting the accuracy after the personalization that is personalized and boost the HAR accuracy when dealing with a wide range of target users. ClusterFL [333], proposes a a clustering-based fl approach for HAR training on hte edge. Finally, [334] proposes an On-Device deep learning approach for STM32 microcontrollers by fine-tuning a 1-D CNN model for further HAR personalization.
@@ VIII-A Libraries and Frameworks for edge learning in edge devices
The most popular frameworks for training deep learning models, PyTorch [349] and TensorFlow [350] have both developed libraries for edge learning. However, while TensorFlow lite [346], allow for both the inference and training of DL models at the edge, PyTorch mobile [351] and ExecuTorch [352], edge learning libraries for PyTorch at the edge, only allow for inference at the time of writing this article. However, it is possible to train PyTorch models using the ONNX Runtime [353]. ONNX Runtime is a cross-platform inference and training machine-learning accelerator, with On-Device Training capabilities [345, 354], the tool has deep integration with PyTorch [353], as well as other deep learning libraries such as Hugging Face [355] and TensorFlow to accelerate training and inferencing on multiple platforms including mobile devices (Android, IOS) and supports multiple hardware accelerators and programming languages.
@@
Libraries that support the experimentation and development of FL Algorithms, as well running, training and the deployement of these algorithms in edge devices, include the afformentioned ONNX Runtime that support FL on edge devices as part of its on-device training support [345, 354]. Flower [347] is a library that is aimed toward research and experimentation of FL algorithms, but allow the execution of the algorithms on a variety of edge devices. FedML [348] is a yet another library that allow researcher to experiment with FL algorithms and migrate them to edge devices. PySyft [361] a FL open-source library built as an extension of some DL libraries, such as PyTorch, Keras and Tensorflow, and allow to be run on mobile devices using KotlinSyft  [^3] KotlinSyft: Syft worker for secure on-device machine learning for Android https://github.com/OpenMined/KotlinSyft for Android and SwiftSyft  [^4] SwiftSyft: Syft worker for secure on-device machine learning for iOS https://github.com/OpenMined/SwiftSyft for iOS. And finally, FedERA [362] is a similar library that also includes a verification Module for ensuring the validation of local models to avoid aggregating malicious local models and a carbon emission tracker module to accurately estimate CO2 emissions during the local parameter update phase.
Edge learning frameworks and libraries that are aimed at specific platforms or hardware have also been proposed. They include CoreML [363], an ML inference and training Framework usable in swift for IOS, proposed by Apple to simplify the deployment and training of ML models on IOS environment. And aimed at MCUs, PULP-TrainLib was proposed in [364] to enable On-Device Training for RISC-V Multi-core MCUs. Librairies and frameworks for other approaches then FL include SplitEasy [101] a framework for training split learning algorithms on mobile devices. And EdgeRL [246] for On-Device Reinforcement Learning, a Light-Weight C/C++ Framework, designed to run on a single-core processor that is typically included in a resource-limited embedded platform.
@@ VIII-B Simulation tools and emulators for edge learning
Emulators and simulation tools are both important in the context of edge computing. simulation tools are used to model the behaviour of fog/edge infrastructures, while emulators provide a way to emulate realistic functionality and configurations, allowing for the study of interoperability across different layers and protocols in edge-cloud environments [365]. In the context of EL, simulators and emulators, allow for the experimentation of ML models on cloud servers or a single device, while emulating the resources, memory and energy constraints of real edge devices enabling rapid prototyping and experimentation.
Most simulation tools available in the context of edge learning are aimed at federated learning, Flower [347], FedML [348], PySyft [361] and FedERA [362] that have been presented in the section VIII-A,allow running FL algorithms on edge devices in addition to the simulation. FL_PyTorch [357] is a PyTorch [349] based simulation tool for federated learning, and TensorFlow Federated [356] is a similar tool for the TensorFlow framework [350]. Others Federated learning simulation tools include : LEAF [359]; FedJax [358] a JAX-based open source library; Flute [366] an open source platform with novel optimization, privacy, and communications strategies ; And finally FedLab [360] a lightweight open-source framework that focus on algorithm effectiveness and communication efficiency, and allows customization on server optimization, client optimization, communication agreement, and communication compression.
@@ IX-I Frameworks to implement training
Despite the proliferation of frameworks and libraries aimed at enabling ML on edge devices, the current landscape lacks robust support for on-device training. Existing tools, such as PyTorch Mobile [351] and ExecuTorch [352], predominantly emphasize inference on mobile/edge devices, neglecting the essential backpropagation algorithms crucial for the training phase. This imbalance in focus between training and inference highlights a critical gap in the current ecosystem. Although some tools have been proposed to facilitate on-device training at the edge VIII-A, there is a pressing need for the development of new libraries, frameworks, and tools explicitly designed for edge learning. Foreseeing the imminent surge in interest, we anticipate an influx of innovative solutions in the coming years, addressing the scarcity of options for on-device ML training. Furthermore, this surge in interest is likely to catalyse advancements in Internet of Things (IoT) and mobile hardware, optimizing their design and capabilities for ML training on the edge.

### 13 Mar 2024 v2

@@ VI-B Reinforcement learning
Reinforcement learning (RL) has been successfully applied in the past on different problems in areas such as robotics, recommendation systems, video games and automatic vehicles [226, 227], making RL a promising and interesting direction for edge learning. However, The training of RL models in resources constrained environments is often limited by high compute and memory requirements from gradient computations [228], making the application of RL in the context of edge learning challenging. Despite these challenges, multiple RL approaches have been proposed for training ML models in the edge. Among them, Federated Reinforcement Learning (FRL) [229, 230] is a promising approach that allow multiple RL agents to learn optimal control policies for a series of devices with slightly different dynamics [231]. FRL is employed to achieve diverse objectives including personalization [232, 233], IoT traffic management [234, 235], Autonomous Systems [231] and resource allocation for unmanned aerial vehicle (UAV) [236]. FedMC [95] integrates RL models trained by multiple edge devices into a general model based on a meta-learning approach. FedGPO is an RL-based aggregation technique for FL introduced in [54], and aims to optimize the energy-efficiency of FL while guaranteeing model convergence. On the other hand, [237] introduces DDQN-Trust, a trust-based double deep Q-learning-based selection algorithm for FL that takes into account the trust scores and energy levels of the IoT devices to make appropriate scheduling decisions and integrate it with the main FL aggregation techniques (FedAvg, FedProx, FedShare and FedSGD). Other Federated Reinforcement Learning methods include [75, 238, 239, 240, 241, 242]
@@ VI-C Semi-supervised learning
As mentioned in the section VI-A using unlabeled data to train on the edge is very promising considering the vast amount of unlabeled data in edge and end devices [32]. Semi-supervised learning, constituting a machine learning paradigm, seeks to leverage both labeled and unlabeled data for specific learning tasks [249], presenting a means to harness the vast amount of unlabeled data at the edge, circumventing the challenges associated with supervised and unsupervised learning. Semi-supervised learning also provides other advantage on the edge, as training datasets are often not completely defined before training and might need to be completed using real-time data [250]. Naturally, several methodologies have surfaced to adapt semi-supervised learning on the edge, using federated learning based approaches [251, 252], or other learning techniques [253, 254, 255, 211, 256, 250, 250, 257].
@@ VII-B Smart Technologies
Smart cities are urban ecosystems designed using IoT technologies to solve urban life problems and improve the residents’ quality of life [279]. In this context, edge learning has been proposed to solve different challenges, mainly for its ability to leverage ML capabilities while preserving network bandwidth and reducing the charge on cloud servers. In [280], a cloud-aided edge learning based on knowledge fusion for smart lighting system has been proposed. Another application of ML in the edge is in smart grid systems where ML is needed to improve demand forecasting and automated demand response, as well as to analyse data related to energy use and obtain energy consumption patterns [281], detect anomalies [282], improve communications [283] and security [284] in the system. Other applications in smart cities include the detection of abnormal and dangerous activities [285, 286], pedestrian detection [287], water consumption forcasting [42, 288] and reducing congestions in intelligent traffic systems [289, 290]. Other more general EL approaches for smart cities include [291, 292, 293, 276, 252].
@@ VII-C Autonomous vehicles
Autonomous vehicles are vehicles that can operate without human intervention, they utilize sensor technologies, AI, and networking to navigate and make decisions [295]. Autonomous vehicles include self-driving cars, trucks, buses, drones, unmanned aerial vehicles (UAVs), and even small robots. As demonstrated in [296] offloading deep learning tasks to edge devices or servers can improve the inference accuracy while meeting the latency constraint, which makes edge learning prefectly suitable for this use-cases, and as expected there has been extensive research done in this area.
Unmanned Aerial Vehicles (UAVs) are by far the most prominent use of EL for autonomous vehicles. In [297], a Synchronous Federated Learning (SFL) structure for multi-UAVs was proposed, that aims to resolve device privacy concerns that comes from sending raw data to UAV servers, as well as UAVs’ limited processing or communication resources. On the other hand, [298] propose a model-aided federated MARL algorithm to coordinate multiple UAVs on data harvesting missions with limited knowledge about the environment, significantly reducing the real-world training data demand. As mentioned previously in section VII-B, [277] aims to assist farming opperations using UAVs that measure soil moisture and ambient environmental conditions and [299] proposes a model to derive computation specifications for learning-based visual odometry (VO) from physical characteristics of UAVs. Other EL applications for UAVs include [236, 300, 301, 302, 303, 299, 304, 305, 306, 307]
@@ VII-D Recommendation systems and personalization
Personalized crowdsourced livecast are another part of personalization methods that might benefit from being offloded to the edge. In [327] the rapid development of crowdsourced livecast and the challenges in providing personalized Quality of Experience (QoE) to viewers is discussed, and it introduces an intelligent edge-learning-based framework called ELCast, which integrates Convolutional Neural Networks (CNN) and deep reinforcement learning (DRL) in edge computing architectures for personalized crowdcast recommendation. In the area related to video games, personalisation involves constructing a system capable of adapting video game rules and content to better suit some aspect of the player preferences, personality, experience and performances [328]. Although not yet explored in edge and on-device learning, [329] propose a Deep Q network model to personalize games based on user-interaction on the edge.
@@ VII-E Human Activity Recognition
Human activity recognition (HAR) refers to the automation of the identification and categorization of the various activities performed by humans and their interactions with the environment [330]. Since personalization for HAR has been shown to improve the results and performances of these systems [331], training the model on the edge using Incremental learning or meta learning approaches can help achieve that while increasing the privacy and reducing the bandwidth consumption. PILOTE [127] proposes an incremental learning based approach for HAR that is intended to be trained on edge devices with extremely limited resources and provides reliable performance in handling catastrophic forgetting. While [332] proposes an approach that proposes a personalizable lightweight CNN model for HAR. As well as a training algorithm to find personalization-friendly parameters, for boosting the accuracy after the personalization that is personalized and boost the HAR accuracy when dealing with a wide range of target users. ClusterFL [333], proposes a clustering-based FL approach for HAR training on the edge. Finally, [334] proposes an On-Device deep learning approach for STM32 microcontrollers by fine-tuning a 1-D CNN model for further HAR personalization.
@@ VIII-A Libraries and Frameworks for edge learning in edge devices
The most popular frameworks for training deep learning models, PyTorch [349] and TensorFlow [350] have both developed libraries for edge learning. However, while TensorFlow lite [346], allows for both the inference and training of DL models at the edge, PyTorch mobile [351] and ExecuTorch [352], Edge ML libraries for PyTorch at the edge, only allow for inference at the time of writing this article. However, it is possible to train PyTorch models using the ONNX Runtime [353]. ONNX Runtime is a cross-platform inference and training machine-learning accelerator, with On-Device Training capabilities [345, 354], the tool has deep integration with PyTorch [353], as well as other deep learning libraries such as Hugging Face [355] and TensorFlow to accelerate training and inferencing on multiple platforms including mobile devices (Android, IOS) and supports multiple hardware accelerators and programming languages.
@@
Libraries that support the experimentation and development of FL Algorithms, as well running, training and the deployement of these algorithms in edge devices, include the afformentioned ONNX Runtime that allows for FL on edge devices as part of its on-device training support [345, 354]. Flower [347] is a library that is aimed toward research and experimentation of FL algorithms, but allows the execution of the algorithms on a variety of edge devices. FedML [348] is a yet another library that allow researchers to experiment with FL algorithms and migrate them to edge devices. PySyft [361] an FL open-source library built as an extension of some DL libraries, such as PyTorch, Keras and Tensorflow, and can to be run on mobile devices using KotlinSyft  [^3] KotlinSyft: Syft worker for secure on-device machine learning for Android https://github.com/OpenMined/KotlinSyft for Android and SwiftSyft  [^4] SwiftSyft: Syft worker for secure on-device machine learning for iOS https://github.com/OpenMined/SwiftSyft for iOS. And finally, FedERA [362] is a similar library that also includes a verification module for ensuring the validation of local models to avoid aggregating malicious local models and a carbon emission tracker module to accurately estimate CO2 emissions during the local parameter update phase.
Edge learning frameworks and libraries that are aimed at specific platforms or hardware have also been proposed. They include CoreML [363], an ML inference and training framework usable in Swift for iOS, proposed by Apple to simplify the deployment and training of ML models on IOS environment. Aimed at MCUs, PULP-TrainLib was proposed in [364] to enable on-device Training for RISC-V Multi-core MCUs. Librairies and frameworks for approaches other than FL include SplitEasy [101] a framework for training split learning algorithms on mobile devices. And EdgeRL [246] for on-device reinforcement learning, a Light-Weight C/C++ Framework, designed to run on a single-core processor that is typically included in a resource-limited embedded platform.
@@ VIII-B Simulation tools and emulators for edge learning
Emulators and simulation tools are both important in the context of edge computing. simulation tools are used to model the behaviour of fog/edge infrastructures, while emulators provide a way to emulate realistic functionalites and configurations, allowing for the study of interoperability across different layers and protocols in edge-cloud environments [365]. In the context of EL, simulators and emulators, allow for the experimentation of ML models on cloud servers or a single device, while emulating the resources, memory and energy constraints of real edge devices enabling rapid prototyping and experimentation.
Most simulation tools available in the context of edge learning are aimed at federated learning, Flower [347], FedML [348], PySyft [361] and FedERA [362] that have been presented in the section VIII-A,allow running FL algorithms on edge devices in addition to the simulation. FL_PyTorch [357] is a PyTorch [349] based simulation tool for federated learning, and TensorFlow Federated [356] is a similar tool for the TensorFlow framework [350]. Other federated learning simulation tools include : LEAF [359]; FedJax [358] a JAX-based open source library; Flute [366] an open source platform with novel optimization, privacy, and communication strategies ; And finally FedLab [360] a lightweight open-source framework that focus on algorithm effectiveness and communication efficiency, and allows customization on server optimization, client optimization, communication agreement, and communication compression.
@@ IX-I Frameworks to implement training
Despite the proliferation of frameworks and libraries aimed at enabling ML on edge devices, the current landscape lacks robust support for on-device training. Existing tools, such as PyTorch Mobile [351] and ExecuTorch [352], predominantly emphasize inference on mobile/edge devices, neglecting the essential backpropagation algorithms crucial for the training phase. This imbalance in focus between training and inference highlights a critical gap in the current ecosystem. Although some tools have been proposed to facilitate on-device training at the edge VIII-A, there is a pressing need for the development of new libraries, frameworks, and tools explicitly designed for edge learning. Foreseeing the surge in interest, we anticipate an influx of innovative solutions in the coming years, addressing the scarcity of options for on-device ML training. Furthermore, this surge in interest is likely to catalyse advancements in IoT and mobile hardware, optimizing their design and capabilities for ML training on the edge.
@@ X Conclusion
Acknowledgments
This research is supported by the Technology Innovation Institute, UAE under the research contract number TII/DSRC/2022/3143.